OpenStudy (hang254):

Find an explicit rule for the nth term of the sequence. -4, -8, -16, -32, ...

OpenStudy (hang254):

@UnkleRhaukus @aaronq

OpenStudy (hang254):

@ParthKohli

ganeshie8 (ganeshie8):

see any pattern ?

OpenStudy (hang254):

it multiplies by 2 after each number in the sequence

ganeshie8 (ganeshie8):

exactly ! so can you write an explicit rule for getting "nth" term ?

OpenStudy (hang254):

2 x -4^n ?

ganeshie8 (ganeshie8):

1st term = -4*2^0 2nd term = -4*2^1 3rd term = -4*2^2 ... nth term = ?

OpenStudy (hang254):

an = -4*2^n

ganeshie8 (ganeshie8):

Very close !! plugin n =1, do u really get 1st term ?

OpenStudy (hang254):

oh, no an = -4*2^n-1 ?

ganeshie8 (ganeshie8):

Yes :)

OpenStudy (hang254):

thanks ganshie!

ganeshie8 (ganeshie8):

\[\large a_n = -4*2^{n-1}\]